# Motion2VecSets: 4D Latent Vector Set Diffusion for Non-rigid Shape Reconstruction and Tracking

### 12 Jan 2024 v1

@@ Abstract
†† Equal Contribution.We introduce Motion2VecSets, a 4D diffusion model for dynamic surface reconstruction from point cloud sequences. While existing state-of-the-art methods have demonstrated success in reconstructing non-rigid objects using neural field representations, conventional feed-forward networks encounter challenges with ambiguous observations from noisy, partial, or sparse point clouds. To address these challenges, we introduce a diffusion model that explicitly learns the shape and motion distribution of non-rigid objects through an iterative denoising process of compressed latent representations. The diffusion-based prior enables more plausible and probabilistic reconstructions when handling ambiguous inputs. We parameterize 4D dynamics with latent vector sets instead of using a global latent. This novel 4D representation allows us to learn local surface shape and deformation patterns, leading to more accurate non-linear motion capture and significantly improving generalizability to unseen motions and identities. For more temporal-coherent object tracking, we synchronously denoise deformation latent sets and exchange information across multiple frames. To avoid the computational overhead, we design an interleaved space and time attention block to alternately aggregate deformation latents along spatial and temporal domains. Extensive comparisons against the state-of-the-art methods demonstrate the superiority of our Motion2VecSets in 4D reconstruction from various imperfect observations, notably achieving a 19% improvement in Intersection over Union (IoU) compared to CaDex [26] for reconstructing unseen individuals from sparse point clouds on the DeformingThings4D-Animals dataset [29].
![[Uncaptioned image]](extracted/5344229/fig/0_teaser/teaser_x2.jpg)
@@ 1 Introduction
Our world, dynamic in its 4D nature, demands an increasingly sophisticated understanding and simulation of our living environment. This offers significant potential for practical applications, including Virtual Reality (VR), Augmented Reality (AR), and robotic simulations.
@@
There have been notable advancements in 3D object modeling, particularly in representations through parametric models [31, 37, 46, 28, 65]. Unfortunately, these template-based models are not suitable to effectively capture the 4D dynamics of general non-rigid objects, due to the assumption of a fixed template mesh. The model-free approaches [33, 52, 26] represent a significant advancement by using coordinate-MLP representations for deformable object reconstruction with arbitrary topologies and non-unified structures. However, these state-of-the-art methods still encounter challenges when facing ambiguous observations of noisy, sparse, or partial point clouds since it is an ill-posed problem where multiple possible reconstructions can match with the input. Besides, they represent dynamics as a sequence of single latent codes and thus struggle to capture shape and motion priors accurately. These issues become even more severe with unseen identities, due to the limited generalizability of global latent representation.
To address above mentioned challenges, we propose Motion2VecSets, a diffusion model designed for 4D dynamic surface reconstruction from sparse, noisy, or partial point clouds.
It explicitly learns the joint distribution of non-rigid object surfaces and temporal dynamics through an iterative denoising process of compressed latent representations. This enables more realistic and varied reconstructions, particularly when dealing with ambiguous inputs.
Inspired by the observation that objects with varying topologies often share similar local geometry and deformation patterns, we represent dynamic surfaces as a sequence of latent vector sets to preserve local shape and deformation details.
One for shape modeling of the initial frame and others to describe the temporal evolution from the initial frame. This latent vector set representation naturally enables the learning of more accurate shape and motion priors, enhancing the model’s generalization capacity to unseen identities and motions.
@@
Specifically, we introduce Synchronized Deformation Vector Set Diffusion, which simultaneously denoises the deformation latent sets across different time frames to enforce spatio-temporal consistency over dynamic surfaces. To manage the memory consumption associated with multiple deformation latent set diffusions, we design interleaved space and time attention block as the basic unit for the denoiser. These blocks aggregate deformation latent sets along spatial and temporal domains in an alternating manner. As illustrated in Figure 1, our Motion2VecSets can reconstruct more plausible non-rigid object surfaces with complicated structures and achieve more robust motion tracking than the existing state-of-the-art.
Our contributions can be summarized as follows:
@@
We introduce a 4D neural representation with latent vector sets, significantly enhancing the capacity to represent complicated shapes as well as motions and improving generalizability to unseen identities and motions.
@@
Extensive comparisons against the state-of-the-art methods demonstrate the superiority of our Motion2VecSets in dynamic surface reconstruction on the Dynamic FAUST [4] and the DeformingThings4D-Animals [29] datasets.
@@ 2 Related works
Traditional methods in 3D representation have primarily employed meshes [59, 19, 30, 49, 39, 50], point clouds [16, 1, 63], and voxels [20, 13, 17, 48, 44] to depict geometry. Complementing these are parametric models, which have effectively represented specific shape categories, such as human bodies (e.g., SMPL [31], STAR [37]), faces (e.g., FLAME [28]), hands (e.g., MANO [46]), and animals (e.g., SMAL [65]). However, these parametric approaches often rely on fixed template topologies, which can lead to difficulties in accurately modeling general non-rigid objects without consistent topological structures. Meanwhile, recent advancements in 3D representation are increasingly adopting implicit methods [10, 18, 33, 40, 11, 6, 58, 61, 7, 51], known for their greater flexibility to represent objects with arbitrary topologies. Notably, Occupancy Networks [33] and DeepSDF [40] employ a continuous implicit framework, enabling the representation of volumetric grids with potentially infinite resolution.
@@
Recent advancements have successfully extended 3D representations to 4D, which captures object dynamics more effectively [38, 36, 5, 24, 52, 26, 16, 54, 27]. For instance, OFlow [36] incorporates Neural-ODE [56] to simulate deformations. LPDC [52] replaces Neural-ODE with an MLP and learns local spatio-temporal codes, capturing both shape and deformations. CaDeX [26] employs a learnable deformation embedding between each frame and its canonical shape. Despite these advancements, methods relying on either ODE [36] solvers or a single global latent vector [52, 26, 38] coupled with an MLP network face challenges in accurately capturing complex real-world 4D dynamics, particularly in non-rigid objects. Drawing inspiration from 3DShape2Vecset [61], which uses a set of latent codes to represent similar local geometry patterns across objects, our proposed method leverages a similar approach to characterize 4D dynamics. Different objects share similar local deformation patterns, our framework uniquely assigns a distinct learnable latent code to each local region, significantly enhancing its ability to precisely model and generalize to unseen identities and motions.
@@
Diffusion models [21], known for their Markov chain-based denoising capability, have gained impressive progress in multiple domains, including image and video processing [8, 35, 15, 22, 34], 3D vision [32, 45, 3, 12, 23, 57, 55, 53]. These models are adept at capturing complex data distributions. In the field of 3D vision, their applications are varied: Luo et al. [32] have successfully applied diffusion models to point cloud generation, and Rombach et al. [45, 3] have adapted them for latent space representations. Additionally, integration with PointNet [42] and triplane features [41], as seen in DiffusionSDF [12], has further enhanced their training capabilities. Concurrent work NAP [27] advances 3D object generation by effectively modeling articulated objects with a novel parameterization and diffusion-denoising approach. A key challenge in representing 4D dynamics with existing diffusion models is their tendency to adapt 3D models directly to 4D and process each frame independently, which can result in discontinuities in temporal and spatial relationships. To bridge this gap, our approach implements a synchronous denoising process for sets of codes. This innovation ensures not only a reduction in spatial complexity but also consistent deformations in latent space. Moreover, recent works [23, 57, 14, 60, 47, 43, 62] in the field of 3D pose estimation and generation also indicate the power of diffusion models. DiffPose [23] utilizes the diffusion model to handle very ambiguous poses and can even predict an infinite number of poses. PhysDiff [60] produces physically plausible motions by incorporating a physics-based motion projection within its diffusion process. However, These methods are still in the realm of pose, our method expands the application of diffusion models to a broader range of deformable surfaces of general non-rigid objects.
@@ 3 Approach
The inputs are frames of sparse or partial noisy point clouds, represented by , where , represents the number of points. The goal is to reconstruct continuous 3D meshes with high fidelity, denoted as , where and refer to the set of vertices faces of the reconstructed mesh at time frame . Conventional feed-forward models face challenges in handling ambiguous inputs within an ill-posed problem setting. Particularly, when observations are sparse, partial, and noisy, generating meaningful reconstructions becomes highly challenging without robust prior knowledge. To reconstruct high-fidelity dynamic shapes accurately, we proposed 4D latent set diffusion to learn shape and motion priors, explicitly learning the distribution of deformable object surface sequences via compressed latent vector sets. While the diffusion model enhances realistic surface reconstruction and deformation tracking, generating multi-modal outputs, the latent set representation and transformer architecture provide the capability to capture more accurate geometry and deformation priors.
@@
Previous works often utilize single global codes [36, 52, 26] to represent 4D sequences, potentially losing significant surface geometry and temporal evolution information. To retain as much detail as possible, we advocate the use of two distinct sets of latent vectors. Specifically, the shape latent set is responsible for reconstructing the initial frame, serving as the reference frame, while the deformation correspondences between the reference and subsequent frames are encoded by the deformation latent set. Compared to previous methods [36, 52, 26] relying on a single global code, we assign local latent codes to individual local regions, which significantly improve the network’s capability to accurately model non-linear motions and generalize to unseen identities and motions. Given the fact that different non-rigid objects share similar local geometry and deformation patterns, the latent sets can also increase the generalization ability to handle unseen motions and identities.
@@
Similar to 3DShape2VecSet [61], we utilize a shape autoencoder to compress the surface of the initial frame into a set of latent codes. Concretely, we leverage a transformer encoder to condense the 3D surface of the initial frame into a set of latent vectors denoted as . Here, represents the overall count of codes and denotes their dimensionality. Following that, a cross-attention layer is used to fuse the latent codes for occupancy field prediction through an MLP. Training is processed by minimizing the binary cross-entropy (BCE) loss, which aligns the predicted occupancy with the actual occupancy , refers to the query points:
@@
Consistent with the latent diffusion framework [45], our model incorporates KL-regularization in the latent space to modulate feature diversity. This ensures the preservation of high-level features and keeps coherent global geometric and deformation patterns, which promotes the learning of diffusion models.
In summary, we characterize a sequence through the shape latent set of the initial reference frame, which describes the implicit surface, and deformation latent sets that depict the dense correspondences between the initial reference frame and the subsequent frames.
3.2 4D Latent Vector Set Diffusion
@@ 3.2.2 Synchronized Deformation Diffusion
To adapt those 3D models [37, 28, 46, 65] directly to 4D, the most straightforward approach is frame-by-frame processing, which may lead to potential discontinuities in temporal and spatial correspondence. Another approach is to aggregate all spatial-temporal point clouds, which would significantly increase the time complexity to for a sequence of frames and points each. However, our 4D latent set representation allows us to bypass the need for brute-force attention across spatial and temporal domains. As discussed in Sec. 3.1, the deformation latents at identical spatial positions across different frames correspond to the deformation behaviors of the same local surface region. Leveraging this property, we implement an alternating aggregation approach for the latent features, systematically switching between the spatial and temporal domains. This method not only enhances efficiency but also preserves the accuracy of our model, leading to a reduction in time complexity to in the spatial domain and in the temporal domain.
The details of synchronized deformation diffusion are described as follows. Given sparse input point clouds , we pair subsequent frames with the first reference frame , i.e., . These pairs are encoded into a series of conditional latents via a transformer encoder. Subsequently, these conditional latents, together with the diffused shape latent set in Sec. 3.2.1, are injected into the denoising network as the condition, which provides useful guidance for the network to handle ambiguous scenarios, such as partial observation.
@@
Figure 4 depicts the denoiser network of our proposed synchronized deformation latent set diffusion. The basic unit is the designed Interleave Spatio-temporal Attention Block (ISTA). Each ISTA block contains three attention layers: Space Self-Attention Layer, Conditional Cross-Attention Layer and Time Self-Attention Layer. The Space Self-Attention Layer initiates spatial information exchange within each set of noised deformation codes :
@@ 4 Experiments
To validate our model, we conducted experiments with two datasets. The first, Dynamic FAUST (D-FAUST) [4], focuses on human body dynamics and includes 10 subjects and 129 sequences
with a split of training (70%), validation (10%), and testing (20%) subsets. The second, DeformingThings4D-Animals (DT4D-A) [29], includes 38 identities with a total of 1227 animations, divided into training (75%), validation (7.5%), and test (17.5%) subsets. The training and validation sets are sequences of seen individuals and motions. The test sets are divided into unseen motions and unseen individuals.
Baselines: Our model’s performance in 4D reconstruction is benchmarked against leading methods: OFlow [36], LPDC [52], and CaDex [26]. OFlow employs a Neural-ODE framework [9] to learn shape transformations, representing a distinct approach in 4D modeling. LPDC, in contrast, utilizes an MLP-based method for deformation processing, offering a different perspective from OFlow. CaDex introduces a unique approach by employing a learned canonical shape for improved stability, setting a novel standard for comparison. To ensure fair evaluation, all methods were tested following the training paradigm established by OFlow [36] for D-FAUST [4], and CaDeX [26] for DT4D-A [29], providing a consistent basis for comparison across the different methodologies.
Evaluation Metrics: We follow [36, 52] and use same metrics. The Intersection over Union (IoU) metric evaluates the overlap between predicted and ground truth meshes; higher IoU values indicate superior accuracy. The Chamfer distance, a bidirectional metric, calculates the average nearest-neighbor distance between two point sets. Lower chamfer distance values imply a closer approximation to the ground truth. We also utilized the -distance error, which measures the Euclidean distance between corresponding points on the predicted and ground truth meshes, with lower values reflecting higher fidelity.
Implementations: In our two-stage training approach, the first stage involves a shape and deformation auto-encoder. The input point cloud with points is randomly sampled from the object surface and the near-surface region. We trained the auto-encoders on two NVIDIA RTX 4090 GPUs with 24 GB memory each, using a batch size of 24 for 100 epochs. The learning rate is set to , with KL-divergence loss weights at (shape) and (deformation). We choose the number of latent codes with channels.
The second stage is the shape and deformation diffusion, we leverage surface points and the encoder we obtained in the first stage to get the target latent codes for diffusion models. We maintained the same learning rate and trained for 50 epochs.
@@
We initially assessed our models’ ability for 4D reconstruction from sparse and noisy point cloud sequences Consistent with the setup in OFlow [36], our network processed sequences of continuous frames. Each frame represented a sparse point cloud with a size of for D-FAUST [4] or for DT4D-A [29]. We also simulate noisy observations by augmenting them with Gaussian noise with a standard deviation of 0.05.
Quantitatively, our model demonstrates superior performance over state-of-the-art models on the D-FAUST [4] and DT4D-A [29] datasets, as detailed in Table 2. This superiority is particularly notable in the unseen individual category of the DT4D-A dataset, which features more diverse topologies from various animals. Here, our model achieves an IoU that is 19% higher than the preceding state-of-the-art model. Additionally, both the chamfer distance and the -correspondence error are reduced to less than half of those recorded by the previous state-of-the-art methods.
Qualitatively, as illustrated in Figure 6 through chamfer distance error maps, our model outperforms in reconstructing complete shapes and minimizing chamfer distance errors, particularly in capturing fast-moving structures like feet of humans and heads of animals.
The superiority of our model is attributed to the proposed 4D latent vector set diffusion, enabling more precise capture of local geometry and deformation patterns. Methods like LPDC [52] and OFlow [36] perform well in human settings thanks to similar human topologies, while CaDex [26] benefits from canonical shape learning. However, the diverse topologies and scales in animal setup, such as between species like foxes and dragons, present a significant challenge for models that optimize global codes. Our approach, in contrast, effectively captures these complex 4D dynamics of general non-rigid objects.

### 13 Apr 2024 v2

@@ Abstract
We introduce Motion2VecSets, a 4D diffusion model for dynamic surface reconstruction from point cloud sequences. While existing state-of-the-art methods have demonstrated success in reconstructing non-rigid objects using neural field representations, conventional feed-forward networks encounter challenges with ambiguous observations from noisy, partial, or sparse point clouds. To address these challenges, we introduce a diffusion model that explicitly learns the shape and motion distribution of non-rigid objects through an iterative denoising process of compressed latent representations. The diffusion-based priors enable more plausible and probabilistic reconstructions when handling ambiguous inputs. We parameterize 4D dynamics with latent sets instead of using global latent codes. This novel 4D representation allows us to learn local shape and deformation patterns, leading to more accurate non-linear motion capture and significantly improving generalizability to unseen motions and identities. For more temporally-coherent object tracking, we synchronously denoise deformation latent sets and exchange information across multiple frames. To avoid computational overhead, we designed an interleaved space and time attention block to alternately aggregate deformation latents along spatial and temporal domains. Extensive comparisons against state-of-the-art methods demonstrate the superiority of our Motion2VecSets in 4D reconstruction from various imperfect observations.
![[Uncaptioned image]](extracted/5532480/fig/0_teaser/teaser_x2.jpg)
@@ 1 Introduction
Our world, dynamic in its 4D nature, demands an increasingly sophisticated understanding and simulation of our living environment. This offers significant potential for practical applications, including Virtual Reality (VR), Augmented Reality (AR), and robotic simulations. There have been notable advances in 3D object modeling, particularly in representations through parametric models [31, 37, 46, 28, 69]. Unfortunately, these template-based models are not effectively suited to capture the 4D dynamics of general non-rigid objects, due to the assumption of a fixed template mesh. Model-free approaches [33, 53, 26] represent a significant advance by using coordinate-MLP representations for deformable object reconstruction with arbitrary topologies and non-unified structures. However, these state-of-the-art methods still encounter challenges when facing ambiguous observations of noisy, sparse, or partial point clouds since it is an ill-posed problem where multiple possible reconstructions can match the input. In addition, they represent dynamics as a sequence of single latent codes and thus struggle to capture shape and motion priors accurately. These issues become even more severe with unseen identities due to the limited generalizability of global latent representation.
@@
To address the above-mentioned challenges, we propose Motion2VecSets, a diffusion model designed for 4D dynamic surface reconstruction from sparse, noisy, or partial point clouds. It explicitly learns the joint distribution of non-rigid object surfaces and temporal dynamics through an iterative denoising process of compressed latent representations. This enables more realistic and varied reconstructions, particularly when dealing with ambiguous inputs. Inspired by the observation that objects with varying topologies often share similar local geometry and deformation patterns, we represent dynamic surfaces as a sequence of latent sets to preserve local shape and deformation details: one for shape modeling of the initial frame and others for describing the temporal evolution from the initial frame. This latent set representation naturally enables the learning of more accurate shape and motion priors, enhancing the model’s generalization capacity to unseen identities and motions. Specifically, we introduce the Synchronized Deformation Vector Set Diffusion, which simultaneously denoises the deformation latent sets across different time frames to enforce spatio-temporal consistency over dynamic surfaces. To manage the memory consumption associated with multiple deformation latent set diffusions, we design an interleaved space and time attention block as the basic unit for the denoiser. These blocks aggregate deformation latent sets along spatial and temporal domains alternately. As illustrated in Figure 1, our Motion2VecSets can reconstruct more plausible non-rigid object surfaces with complicated structures and achieve more robust motion tracking than the state-of-the-art method. Our contributions can be summarized as follows:
@@
We introduce a 4D neural representation with latent sets, significantly enhancing the capacity to represent complicated shapes as well as motions and improving generalizability to unseen identities and motions.
@@
Extensive comparisons against state-of-the-art methods demonstrate the superiority of our Motion2VecSets in dynamic surface reconstruction on the Dynamic FAUST [4] and the DeformingThings4D-Animals [29] datasets.
@@ 2 Related works
Traditional methods in 3D representation have primarily used meshes [60, 19, 30, 50, 39, 51], point clouds [16, 1, 66], and voxels [20, 13, 17, 49, 44] to represent geometry. Complementing these are parametric models, which have effectively modeled specific shape categories, such as human bodies (e.g., SMPL [31], STAR [37]), faces (e.g., FLAME [28]), hands (e.g., MANO [46]), and animals (e.g., SMAL [69]). However, these parametric approaches often rely on fixed templates, which can result in difficulties accurately modeling general non-rigid objects without consistent topological structures. Meanwhile, recent advancements in 3D representation are increasingly using implicit methods [10, 18, 33, 40, 11, 6, 59, 64, 7, 52, 63], known for their greater flexibility to represent objects with arbitrary topologies. In particular, Occupancy Networks [33] and DeepSDF [40] employ a continuous implicit framework, enabling the representation of volumetric grids offering potentially infinite resolution.
@@
Recent advancements have successfully extended 3D representations to 4D, which more effectively captures object dynamics [38, 36, 5, 24, 53, 26, 16, 54, 27, 48, 68]. For example, OFlow [36] incorporates Neural-ODE [57] for simulating deformations. LPDC [53] replaces Neural-ODE with an MLP and learns local spatio-temporal codes, capturing both shape and deformations. CaDeX [26] employs a learnable deformation embedding between each frame and its canonical shape. However, methods relying on either ODE [36] solvers or a single global latent vector [53, 26, 38] coupled with an MLP network face challenges in capturing complex real-world 4D dynamics, particularly in non-rigid objects. Drawing inspiration from 3DShape2Vecset [64], which uses a set of latent codes to represent similar local geometry patterns across objects, our proposed method leveragesa similar approach for characterizing 4D dynamics. Different objects share similar local deformation patterns, our framework uniquely assigns a distinct learnable latent code to each local region, significantly enhancing their ability to precisely model and generalize to unseen identities and motions.
@@
Diffusion models [21], known for their Markov chain-based denoising capability, have made impressive progress in multiple tasks, including image and video processing [8, 35, 15, 22, 34], 3D vision [32, 45, 3, 12, 23, 58, 56, 55, 62]. These models are adept at capturing complex data distributions. In the field of 3D vision, their applications are varied: Luo et al. [32] have successfully applied diffusion models to point cloud generation, and Rombach et al. [45, 3] have adapted them for latent space representations. Additionally, integration with PointNet [42] and triplane features [41], as seen in DiffusionSDF [12], has further enhanced their training capabilities. Concurrent work NAP [27] advances 3D object generation by effectively modeling articulated objects with a novel parameterization and diffusion-denoising approach. A key challenge in representing 4D dynamics with existing diffusion models is their tendency to adapt 3D models directly to 4D and process each frame independently, which can result in discontinuities in temporal and spatial relationships. To bridge this gap, our approach implements synchronous denoising processes for sets of codes. This innovation ensures not only a reduction in spatial complexity but also consistent deformations in latent space. Moreover, recent works [23, 58, 14, 61, 47, 43, 65] in the field of 3D pose estimation and generation also indicate the power of diffusion models. DiffPose [23] utilizes the diffusion model to handle very ambiguous poses and can even predict an infinite number of poses. PhysDiff [61] produces physically plausible motions by incorporating a physics-based motion projection within its diffusion process. However, these methods are still in the realm of pose, our method expands the application of diffusion models to a broader range of deformable surfaces of general non-rigid objects. Similar to the concurrent work DPHMs [55], our approach utilizes diffusion priors to facilitate robust 4D reconstruction from imperfect observations.
@@ 3 Approach
The inputs are frames of sparse, partial, or noisy point clouds, represented by , where , represents the number of points. The goal is to reconstruct continuous 3D meshes with high fidelity, denoted as , where and refer to the set of vertices faces of the reconstructed mesh at time frame . Conventional feed-forward models face challenges in handling ambiguous inputs within an ill-posed problem setting. Particularly, when observations are sparse, partial, and noisy, generating meaningful reconstructions becomes highly challenging without robust prior knowledge. To reconstruct high-fidelity dynamic shapes accurately, we proposed 4D latent set diffusion to learn shape and motion priors, explicitly learning the distribution of deformable object surface sequences via compressed latent vector sets. While the diffusion model enhances realistic surface reconstruction and deformation tracking, generating multi-modal outputs, the latent set representation and transformer architecture provide the capability to capture more accurate geometry and deformation priors.
@@
Previous works often utilize single global codes [36, 53, 26] to represent 4D sequences, potentially losing significant surface geometry and temporal evolution information. To retain as much detail as possible, we advocate the use of two distinct sets of latent vectors. Specifically, the shape latent set is responsible for reconstructing the initial frame, serving as the reference frame, while the deformation correspondences between the reference and subsequent frames are encoded by the deformation latent set. Compared with previous methods [36, 53, 26] relying on a single global code, we assign local latent codes to individual local regions, which significantly improves the network’s capability to accurately model non-linear motions and generalize to unseen identities and motions. Given that different non-rigid objects share similar local geometry and deformation patterns, the latent sets can also increase the generalization ability to handle unseen motions and identities.
@@
Similar to 3DShape2VecSet [64], we utilize a shape autoencoder to compress the surface of the initial frame into a set of latent codes. Concretely, we leverage a transformer encoder that condenses the 3D surface of the initial frame into a set of latent vectors denoted as . Here, represents the overall count of codes and denotes their dimensionality. Following that, a cross-attention layer is used to fuse the latent codes for occupancy field prediction through an MLP. Training involves minimizing the binary cross-entropy (BCE) loss, which aligns the predicted occupancy with the actual occupancy , refers to the query points:
@@
Consistent with the latent diffusion framework [45], our model incorporates KL-regularization in the latent space to modulate feature diversity. This ensures the preservation of high-level features and keeps coherent global geometric and deformation patterns, which promotes the learning of diffusion models. In summary, we characterize a sequence through the shape latent set of the initial reference frame, which describes the implicit surface, and deformation latent sets that depict the dense correspondences between the initial reference frame and the subsequent frames.
3.2 4D Latent Set Diffusion
@@ 3.2.2 Synchronized Deformation Diffusion
To adapt these 3D models [37, 28, 46, 69] directly to 4D, the most straightforward approach is frame-by-frame processing, which may lead to discontinuities in temporal and spatial correspondence. Another approach is to aggregate all spatial-temporal point clouds, which would significantly increases the time complexity to for a sequence of frames and points each. However, our 4D latent set representation allows us to bypass the need for brute-force attention across spatial and temporal domains. As discussed in Sec. 3.1, the deformation latents at identical spatial positions across different frames correspond to the deformation behaviors of the same local surface region. Leveraging this property, we implement an alternating aggregation approach for the latent features, systematically switching between the spatial and temporal domains. This method not only enhances efficiency but also preserves the accuracy of our model, leading to a reduction in time complexity to in the spatial domain and in the temporal domain. The details of synchronized deformation diffusion are described as follows. Given sparse input point clouds , we pair subsequent frames with the first reference frame , i.e., . These pairs are encoded into a series of conditional latents via a transformer encoder. Then these conditional latents, together with the diffused shape latent set in Sec. 3.2.1, are injected into the denoising network as the condition providing guidance for the network to handle ambiguous scenarios, like partial observation.
@@
Figure 4 depicts the denoiser network of our proposed synchronized deformation latent set diffusion. The basic unit is the designed Interleaved Spatio-temporal Attention Block (ISTA). Each ISTA block contains three attention layers: Space Self-Attention Layer, Conditional Cross-Attention Layer and Time Self-Attention Layer. The Space Self-Attention Layer initiates spatial information exchange within each set of noised deformation codes :
@@ 4 Experiments
We conducted experiments on two 4D datasets. The first, Dynamic FAUST (D-FAUST) [4], focuses on human body dynamics, including 10 subjects and 129 sequences. It is split into training (70%), validation (10%), and test (20%) subsets, following [36]. The second, DeformingThings4D-Animals (DT4D-A) [29], includes 38 identities with a total of 1227 animations, divided into training (75%), validation (7.5%), and test (17.5%) subsets as [26]. The training and validation sets use motion sequences of seen individuals. The test set is divided into two parts: unseen motions and unseen individuals.
Baselines: We compare against state-of-the-art methods in 4D reconstruction, including OFlow [36], LPDC [53], CaDex [26].
OFlow assigns each 4D point both an occupancy value and a motion velocity vector, utilizing a Neural-ODE framework [9] for learning deformations. LPDC employs a MLP to parallelly learn correspondences among occupancy fields across different time steps via explicitly learning continuous displacement vector fields from spatio-temporal shape representation. CaDeX introduces a canonical map factorization and utilizes invertible deformation networks to maintain homeomorphisms.
For fair comparisons, we follow their original training paradigms.
Evaluation Metrics: The Intersection over Union (IoU) evaluates the overlap between predicted and ground truth meshes; The Chamfer distance calculates the average nearest-neighbor distance between two point sets; -distance error measures the Euclidean distance between corresponding points on the predicted and ground truth meshes.
Implementations: The training of our approach consists of two stages. The first stage involves two auto-encoders. The input point clouds () are randomly sampled from object surfaces and near-surface regions. For the shape auto-encoder, the learning rate is , with KL-divergence loss weights . For the deformation auto-encoder, the learning rate is , with KL-divergence loss weights . They are trained for 100 epochs with batch size 24.
The second stage is the diffusion models, the learning rate for both shape and deformation diffusion models is and they are trained for 50 epochs with a batch size of 8 for shape diffusion and 4 for deformation diffusion.
Runtime: The training takes about 60 hours(2*RTX 4090) . The inference takes about 11s for 17 frames(1*RTX 3080).
@@
We initially assessed our models’ ability for 4D reconstruction from sparse and noisy point cloud sequences Consistent with the setup in OFlow [36], our network processed sequences of continuous frames. Each frame represented a sparse point cloud, with for D-FAUST [4] or for DT4D-A [29]. We also simulate noisy observations with Gaussian noise ( = 0.05).
Quantitatively, our model demonstrates superior performance over state-of-the-art models on the D-FAUST [4] and DT4D-A [29] datasets, as detailed in Tab. 2. This superiority is particularly notable in the unseen individual category of the DT4D-A dataset, which features more diverse topologies from various animals. Additionally, both chamfer distance and -correspondence error are reduced to less than half of those recorded by the previous state-of-the-art methods. Qualitatively, as illustrated in Fig. 5, our model outperforms in reconstructing complete shapes and minimizing chamfer distance errors, particularly in capturing fast-moving structures like feet of humans and heads of animals.
The superiority of our model is attributed to the proposed 4D latent set diffusion, enabling a more precise capture of local geometry and deformation patterns. Methods like LPDC [53] and OFlow [36] perform well in human settings thanks to similar human topologies, while CaDex [26] benefits from canonical shape learning. However, the diverse topologies and scales in animal setup, such as dragons, present a significant challenge for models that optimize global codes. Our approach, in contrast, effectively captures these complex 4D dynamics of general non-rigid objects.